OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

Ok, we need to set up two equations. The first one will tell us how many total dimes and quarters you have. The second one will tell us the value of a quarter, the value of a dime, and the total value. Let's have d = # of dimes and q = # of quarters d+q = 46 .10d+.25q=7.00 Are you following me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, now we can use elimination, substitution, or graphing to solve this. See if you can solve it, if you have any trouble let me know.

OpenStudy (anonymous):

Ok so what I've done already is... d=46-q 10(46-q)+25q=7 460-10q+25q=7 460+15q=7 15q=-453 and I got q=-30.2 and I don't think it's right...

OpenStudy (anonymous):

Alright, I think what you did wrong here is when you changed the equation from .10d+.25q=7.00 to 10d+25q=7 you forgot to move the decimal place over two on the 7 aswell. So the second equation should be 10d+25q=700 Now try solving it. :) That was good that you noticed that your problem didn't make sense. If it doesn't make sense, it probably isn't right!

OpenStudy (anonymous):

OHHH I FORGOT THE DECIMALS!!!

OpenStudy (anonymous):

It happens. Those decimals are tricky. :)

OpenStudy (anonymous):

16 quarters?